# OriGen: Enhancing RTL Code Generation with Code-to-Code Augmentation and Self-Reflection

### v1

@@ Abstract.
Recent studies have illuminated that Large Language Models (LLMs) exhibit substantial potential in the realm of RTL (Register Transfer Level) code generation, with notable advancements evidenced by commercial models such as GPT-4 and Claude3-Opus. Despite their proficiency, these commercial LLMs often raise concerns regarding privacy and security. Conversely, open-source LLMs, which offer solutions to these concerns, have inferior performance in RTL code generation tasks to commercial models due to the lack of high-quality open-source RTL datasets.
To address this issue, we introduce OriGen, a fully open-source framework featuring self-reflection capabilities and a dataset augmentation methodology for generating high-quality, large-scale RTL code. We propose a novel code-to-code augmentation methodology that leverages knowledge distillation to enhance the quality of the open-source RTL code datasets. Additionally, OriGen is capable of correcting syntactic errors by leveraging a self-reflection process based on feedback from the compiler. The self-reflection ability of the model is facilitated by a carefully constructed dataset, which comprises a comprehensive collection of samples. Experimental results demonstrate that OriGen remarkably outperforms other open-source alternatives in RTL code generation, surpassing the previous best-performing LLM by 9.8% on the VerilogEval-Human benchmark. Furthermore, OriGen exhibits superior capabilities in self-reflection and error rectification, surpassing GPT-4 by 18.1% on the benchmark designed to evaluate the capability of self-reflection.
@@ 1. Introduction
Recently, large language models (LLMs) have demonstrated outstanding performance in natural language comprehension and generation  (Kalyan et al., 2021). Studies have shown that LLMs exhibit considerable ability in code generation tasks  (Sun et al., 2024). Commercial LLMs, such as GPT-4 (Achiam et al., 2023) and Claude3-Opus (Anthropic, 2024), generate high-quality software code for common programming languages like C++ and Python, which enhance coding productivity significantly. Moreover, LLMs have also illustrated impressive capabilities in hardware code generation, especially for the Register Transfer Level (RTL) code. Specifically, RTL code generation from natural language instructions presents an innovative approach to enhance hardware development productivity, offering the potential to revolutionize existing HDL coding workflows by alleviating the burdensome task of HDL coding for designers.
@@
While commercial LLMs have demonstrated proficiency in generating RTL code, they often raise concerns regarding privacy and security. These issues are especially critical in the domain of hardware design, as RTL code frequently contains valuable intellectual property. Furthermore, the closed-source nature of these LLMs restricts researchers from conducting in-depth analyses and customizations, impeding further fine-tuning of the models in specific domains, such as hardware design. In contrast, open-source models provide better privacy and security while facilitating further improvement and customization. Open-source models, such as StarCoder2 (Lozhkov et al., 2024), DeepSeek-Coder (Guo et al., 2024), and CodeQwen-V1.5 (Bai et al., 2023), have demonstrated promising results in code generation for prevalent programming languages like Python and C/C++, while their performance in RTL code generation still lags behind GPT-3.5. This underscores the pressing need to develop a specialized open-source LLM tailored for RTL code generation.
@@
The scarcity of high-quality RTL code datasets is a significant factor contributing to the poor quality of RTL code generation. In LLM training, both the scale and quality of the dataset are essential for effective code generation performance. However, compared to other popular programming languages, there is a notable shortage of high-quality, large-scale RTL code datasets, which poses a significant challenge for achieving satisfactory results in RTL generation tasks. To address this issue, several studies have focused on collecting open-source code snippets  (Thakur et al., 2023a, b; Pei et al., 2024; Dehaerne et al., 2023). Nevertheless, these open-source datasets may include low-quality code that can impair the model’s performance  (Liu et al., 2023c). Other researchers have developed methodologies for synthesizing RTL code using LLM  (Liu et al., 2023c; Chang et al., 2024; Liu et al., 2023b). VerilogEval (Liu et al., 2023c) generates datasets by adding descriptions to open-source Verilog code but suffers from its average low quality. RTLCoder (Liu et al., 2023b) selects hardware-related keywords as seeds to generate natural language instructions, which are then converted into corresponding RTL code by GPT-3.5. Although this dataset meets quality standards, it is limited in scale because the generated data is restricted to the selected keywords.
@@
Self-reflection is also a critical factor to be considered in the field of RTL code generation. In the context of RTL code generation, self-reflection refers to the model’s assessment of its generated code based on feedback from the compiler and simulator. This process involves identifying the causes of errors and refining the results accordingly, just as in the hardware design process, where RTL code undergoes rigorous evaluation by the compiler and simulator to ensure its correctness and adherence to design specifications.
@@
However, existing open-source LLMs mainly focus on code generation, neglecting the interactive process of compilation and simulation in hardware design methodologies. In consideration of the above limitations, several studies have integrated self-reflection methodologies into their frameworks (Tsai et al., 2023; Thakur et al., 2023c; Wu et al., 2024). Nonetheless, both RTLFixer (Tsai et al., 2023) and AutoChip (Thakur et al., 2023c) rely on commercial LLMs for self-reflection, which raises potential concerns regarding privacy and security, as previously discussed. Open-source LLMs generally exhibit weaker self-reflection capabilities compared to commercial LLMs, primarily due to their limited training data. Consequently, there is a need to develop a methodology for constructing datasets specifically designed to enhance the self-reflection abilities of open-source LLMs.
@@
In this paper, we introduce OriGen, a fully open-source framework featuring self-reflection capabilities and a dataset augmentation methodology. We propose a novel code-to-code augmentation that leverages knowledge distillation to improve the quality of open-source RTL code datasets, which are of a large scale but have limited quality due to the lack of rigorous review and validation processes.
@@
Our approach involves extracting high-level code descriptions from the open-source RTL code and refining the code based on these descriptions using the Claude3-Haiku model, which acts as the teacher model. Commercial LLMs have been trained on vast amounts of high-quality data and possess strong language understanding and generation capabilities. By leveraging these models as teachers, we can distill their knowledge and refine the open-source RTL code. This method enables the generated code to potentially achieve even superior performance of the teacher model.
@@
To evaluate the LLM’s capability of self-reflection in RTL code, we construct a benchmark named VerilogFixEval that comprises 221 cases of failed compilation from VerilogEval (Liu et al., 2023c). OriGen is able to correct syntactic errors by leveraging a self-reflection process based on feedback from the compiler. When the generated code fails to pass the compiler’s verification, the model initiates the self-reflection process. During this phase, the model accepts the erroneous code and compiler error messages to fix the RTL code, enhancing its correctness and reliability. The self-reflection ability of the model is facilitated by a carefully constructed error-correction dataset, which comprises a comprehensive collection of samples. This dataset includes natural language instructions, erroneous code, compiler error messages, and the corrected code generated by Claude3-Haiku. It serves as a valuable resource to bridge the gap in self-reflection capabilities between open-source and commercial closed-source LLMs. By training on this dataset, OriGen can acquire self-reflection abilities comparable to those of commercial LLMs, enabling it to generate more reliable and accurate RTL code.
@@
  • We introduce OriGen, which significantly outperforms other alternatives designed for RTL code generation and achieves performance comparable to the latest version of GPT-4 Turbo.
@@
  • we introduce a self-reflection mechanism that enables OriGen to autonomously fix syntactic errors by leveraging feedback from the compiler, improving its code generation accuracy. Furthermore, we construct a dataset to improve the model’s capability of self-reflection based on compiler error messages and develop a benchmark to evaluate this capability.
Experimental results demonstrate that OriGen significantly outperforms other alternatives designed for RTL code generation  (Pei et al., 2024; Liu et al., 2023b; Nadimi and Zheng, 2024; Liu et al., 2023c, a) surpassing the previous best-performing LLM by 9.8% on the VerilogEval-Human  (Liu et al., 2023c) benchmark. Furthermore, experimental results on VerilogFixEval illustrate that OriGen’s self-reflection capability in RTL code generation not only surpasses other alternatives but also outperforms GPT-4 Turbo by 18.1%.
@@ 2.1. LLMs for Verilog Generation
Both VerilogEval  (Liu et al., 2023c) and RTLCoder  (Liu et al., 2023b) acknowledge the importance of dataset quality and adopt data synthesis methods using commercial LLMs to improve the quality of their datasets. VerilogEval proposes a synthesis approach that utilizes GPT-3.5 to add descriptions to each code snippet in the filtered VeriGen  (Thakur et al., 2023b) dataset. This approach allows the LLM trained on the dataset to learn not only the syntax of Verilog but also to understand the correspondence between Verilog code and natural language. However, due to the average low quality of code in the dataset, VerilogEval still falls short of GPT-3.5. RTLCoder (Liu et al., 2023b) adopts a synthesis approach to generate RTL code from natural language specifications. It selects one or two keywords from a prepared list of hundreds of digital design keywords and utilizes GPT-3.5 to generate RTL design instructions based on these keywords. Subsequently, GPT-3.5 is employed again to generate corresponding RTL code from these instructions. However, the reliance on a limited set of keywords poses challenges for further expanding this dataset, as it may restrict the scale and complexity of the generated RTL code. In contrast, the dataset synthesized through the code-to-code approach is not only of high quality, as demonstrated by the experimental results, but also extensive in scale, encompassing a diverse range of RTL code samples. This approach overcomes the limitations of above synthesis and enables the generation of a comprehensive and high-quality dataset that can effectively support the training of LLMs for RTL code generation.
@@ 2.2. Reflection for Verilog Generation
Recognizing the limitations of previous research on LLMs for RTL generation, AutoChip  (Thakur et al., 2023c) and RTLFixer  (Tsai et al., 2023) decide to introduce self-reflection into the their framework. Specifically, RTLFixer employs GPT to generate and reflect on code. It includes approximately 80 erroneous code snippets that fail compilation, each paired with guidance. During the RTL code generation process, if the generated code fails to compile, RTLFixer  (Tsai et al., 2023) queries the database for the guidance most closely related to the error. However, the generalizability of this approach is limited due to the constraint of the guidance database. In contrast, AutoChip  (Thakur et al., 2023c) directly provides the compiler error messages to the LLM rather than querying a database for guidance.
@@
2.3. Knowledge Distillation
To enable open-source models to acquire capabilities comparable to commercial closed-source models, knowledge distillation has become a focus for researchers.  (Xu et al., 2024). A common approach involves leveraging the closed-source LLM to provide reference answers, which are then used to fine-tune the open-source LLM through Supervised Fine-Tuning (SFT), enabling the latter to learn from the knowledge embodied in the closed-source model. Specifically, SFT is a effective knowledge distillation method. In this approach, the teacher model, typically an advanced LLM, is first used to generate a large number of input-output data pairs. Subsequently, these data pairs generated by the teacher model are utilized as training data to fine-tune the student model, typically a smaller open-source language model. Through this process, the student model can mimic the behavior of the teacher model, thereby acquiring the knowledge within the teacher model.
The code-to-code augmentation methodology employed in this work falls under the realm of knowledge distillation techniques. By leveraging the knowledge within these advanced models, the proposed method can effectively distill and transfer their capabilities to the target model, enabling it to acquire the desired capability of RTL code generation in a scalable and automated manner.
@@ 3.1. Overview
The overview of the code-to-code augmentation methodology and OriGen’s generation and self-reflection framework are illustrated in Figure 1 and Figure 2 respectively. As shown in Figure 1, in the code-to-code augmentation process, two datasets are generated: an enhanced code dataset and an error-correction dataset. The enhanced code dataset contains code-description paired data samples, enabling the LLM to learn the correspondence between Verilog code and natural language. The error-correction dataset comprises erroneous code-fixed code paired data samples, allowing the LLM to enhance its self-reflection capability by learning from examples of code errors and their corrections.
As shown in Figure 2, in the code generation and error correction stage, OriGen comprises a base LLM and two trained LoRA models: Gen LoRA and Fix LoRA. Gen LoRA is trained on the enhanced code dataset and is responsible for the initial generation of RTL code upon receiving natural language instructions. Fix LoRA, which is trained on the error-correction dataset based on Gen LoRA, is tasked with fixing erroneous code that fails to compile. The fix process is implemented as a self-reflection loop, iteratively refining the code until it passes verification or reaches a predetermined maximum number of iterations.
@@ 3.2. Code-to-Code Augmentation
A full example of dataset augmentation is shown in Figure 4.

### v2

@@ Abstract.
Recent studies have demonstrated the significant potential of Large Language Models (LLMs) in generating Register Transfer Level (RTL) code, with notable advancements showcased by commercial models such as GPT-4 and Claude3-Opus. However, these proprietary LLMs often raise concerns regarding privacy and security. While open-source LLMs offer solutions to these concerns, they typically underperform commercial models in RTL code generation tasks, primarily due to the scarcity of high-quality open-source RTL datasets. To address this challenge, we introduce OriGen, a fully open-source framework that incorporates self-reflection capabilities and a novel dataset augmentation methodology for generating high-quality, large-scale RTL code. Our approach employs a code-to-code augmentation technique to enhance the quality of open-source RTL code datasets. Furthermore, OriGen can rectify syntactic errors through a self-reflection process that leverages compiler feedback.
Experimental results demonstrate that OriGen significantly outperforms other open-source alternatives in RTL code generation. It surpasses the previous best-performing open-source LLM by 12.8% and even exceeds GPT-4 Turbo in the pass@1 metric on the VerilogEval-Human benchmark. Moreover, OriGen exhibits superior capabilities in self-reflection and error correction, outperforming GPT-4 by 19.9% on a benchmark designed to evaluate self-reflection capabilities.
OriGen is open source at GitHub(https://github.com/pku-liang/OriGen)
@@ 1. Introduction
Recent advancements in large language models (LLMs) have demonstrated exceptional capabilities in natural language comprehension and generation  (Kalyan et al., 2021). Studies have shown that LLMs exhibit considerable proficiency in code generation tasks  (Sun et al., 2024). Commercial LLMs, such as GPT-4 (Achiam et al., 2023) and Claude3-Opus (Anthropic, 2024), have showcased their ability to generate high-quality software code for common programming languages like C++ and Python, significantly enhancing coding productivity. Moreover, LLMs have demonstrated impressive capabilities in hardware code generation, particularly for Register Transfer Level (RTL) code. The generation of RTL code from natural language instructions presents an innovative approach to enhance hardware development productivity. This method has the potential to revolutionize existing Hardware Description Language (HDL) coding workflows by alleviating the burdensome task of HDL coding for designers. Similar to High-Level Synthesis (HLS), LLMs can streamline the design process by enabling code generation from high-level specifications, thus simplifying complex hardware design tasks (Canis et al., 2011; Coussy et al., 2009).
@@
While commercial LLMs have demonstrated proficiency in generating RTL code, they often raise concerns regarding privacy and security. These issues are particularly critical in the domain of hardware design, as RTL code frequently contains valuable intellectual property. Furthermore, the closed-source nature of these LLMs restricts researchers from conducting in-depth analyses and customizations, impeding further fine-tuning of the models in specific domains, such as hardware design. In contrast, open-source models provide enhanced privacy and security while facilitating further improvement and customization. Open-source models, such as StarCoder2 (Lozhkov et al., 2024), DeepSeek-Coder (Guo et al., 2024), and CodeQwen-V1.5 (Bai et al., 2023), have demonstrated promising results in code generation for prevalent programming languages like Python and C/C++. However, their performance in RTL code generation still falls behind GPT-3.5. This performance gap underscores the pressing need to develop a specialized open-source LLM tailored for RTL code generation.
@@
Nonetheless, there is a notable scarcity of high-quality, large-scale RTL code datasets compared to those available for other popular programming languages. This shortage poses a significant challenge in achieving satisfactory results for RTL generation tasks. To address this issue, several studies have focused on collecting open-source code snippets  (Thakur et al., 2023a, b; Pei et al., 2024; Dehaerne et al., 2023). However, these open-source datasets may include low-quality code that can impair model performance (Liu et al., 2023c). Alternative approaches have been developed to synthesize RTL code using LLMs  (Liu et al., 2023c; Chang et al., 2024; Liu et al., 2023b). VerilogEval (Liu et al., 2023c) generates datasets by adding descriptions to open-source Verilog code but suffers from overall low quality. RTLCoder  (Liu et al., 2023b) selects hardware-related keywords as seeds to generate natural language instructions, which are then converted into corresponding RTL code using GPT-3.5. While this dataset meets quality standards, its scale is limited due to the restriction of generated data to selected keywords.
@@
Self-reflection is another critical factor in RTL code generation. In this context, self-reflection refers to the model’s ability to assess its generated code based on feedback from hardware compilers and simulators. This process involves identifying error causes and refining results accordingly, mirroring the hardware design process where RTL code undergoes rigorous evaluation to ensure correctness and adherence to design specifications.
@@
Existing open-source LLMs primarily focus on code generation, often neglecting the interactive process of compilation and simulation integral to hardware design methodologies. To address these limitations, several studies have integrated self-reflection methodologies into their frameworks  (Tsai et al., 2023; Thakur et al., 2023c; Wu et al., 2024). However, both RTLFixer (Tsai et al., 2023) and AutoChip (Thakur et al., 2023c) rely on commercial LLMs for self-reflection, raising potential privacy and security concerns as previously discussed. Open-source LLMs generally exhibit weaker self-reflection capabilities compared to their commercial counterparts, primarily due to limited training data. Consequently, there is a pressing need to develop methodologies for constructing datasets specifically designed to enhance the self-reflection abilities of open-source LLMs in the context of RTL code generation.
@@
In this paper, we introduce OriGen, a fully open-source framework for RTL generation featuring self-reflection capabilities and a dataset augmentation methodology. We propose a novel code-to-code augmentation technique to improve the quality of open-source RTL code datasets, which are typically large-scale but of limited quality due to the lack of rigorous review and validation processes.
@@
Our approach involves extracting high-level code descriptions from open-source RTL code and refining the code based on these descriptions using the Claude3-Haiku model as a teacher. Commercial LLMs, trained on vast amounts of high-quality data, possess strong programming language understanding and generation capabilities. By leveraging these models as teachers, we can distill their knowledge to refine open-source RTL code, potentially achieving performance superior to the teacher model through rigorous review and validation.
@@
TTo evaluate the LLM’s self-reflection capability in RTL code, we construct a benchmark named VerilogFixEval, comprising 221 cases of failed compilation from VerilogEval (Liu et al., 2023c). OriGen can correct syntactic errors by leveraging a self-reflection process based on compiler feedback. When generated code fails to pass compiler verification, the model initiates the self-reflection process, accepting erroneous code and compiler error messages to fix the RTL code, thereby enhancing its correctness and reliability. The model’s self-reflection ability is facilitated by a carefully constructed error-correction dataset, which includes natural language instructions, erroneous code, compiler error messages, and corrected code generated by Claude3-Haiku. This dataset serves as a valuable resource to bridge the gap in self-reflection capabilities between open-source and commercial closed-source LLMs. By training on this dataset, OriGen can acquire self-reflection abilities comparable to those of commercial LLMs, enabling more reliable and accurate RTL code generation.
@@
  • We introduce OriGen, which significantly outperforms other alternatives designed for RTL code generation and achieves performance comparable to the latest version of GPT-4 Turbo. OriGen is open source at https://github.com/pku-liang/OriGen.
@@
  • We introduce a self-reflection mechanism that enables OriGen to autonomously fix syntactic errors by leveraging compiler feedback, improving its code generation accuracy. Furthermore, we construct a dataset to improve the model’s self-reflection capability based on compiler error messages and develop a benchmark to evaluate this capability.
Experimental results demonstrate that OriGen significantly outperforms other open-source alternatives in RTL code generation (Pei et al., 2024; Liu et al., 2023b; Nadimi and Zheng, 2024; Liu et al., 2023c, a). It surpasses the previous best-performing open-source LLM by 12.8% and even exceeds GPT-4 Turbo in the pass@1 metric on the VerilogEval-Human benchmark. Furthermore, OriGen exhibits superior capabilities in self-reflection and error correction, outperforming GPT-4 by 19.9% on the benchmark designed to evaluate self-reflection capabilities.
@@ 2.1. LLMs for Verilog Generation
Both VerilogEval  (Liu et al., 2023c) and RTLCoder  (Liu et al., 2023b) acknowledge the importance of dataset quality and adopt data synthesis methods using commercial LLMs to improve the quality of their datasets. VerilogEval proposes a synthesis approach that utilizes GPT-3.5 to add descriptions to each code snippet in the filtered VeriGen  (Thakur et al., 2023b) dataset. This approach allows the LLM trained on the dataset to learn not only the syntax of Verilog but also to understand the correspondence between Verilog code and natural language. However, due to the average low quality of code in the dataset, VerilogEval still falls short of GPT-3.5. RTLCoder (Liu et al., 2023b) adopts a synthesis approach to generate RTL code from natural language specifications. It selects one or two keywords from a prepared list of hundreds of digital design keywords and utilizes GPT-3.5 to generate RTL design instructions based on these keywords. Subsequently, GPT-3.5 is employed again to generate corresponding RTL code from these instructions. However, the reliance on a limited set of keywords poses challenges for further expansion of this dataset, as it may restrict the scale and complexity of the generated RTL code. In contrast, the dataset synthesized through the code-to-code approach is not only of high quality, as demonstrated by the experimental results, but also extensive in scale, encompassing a diverse range of RTL code samples. This approach overcomes the limitations of above synthesis and enables the generation of a comprehensive and high-quality dataset that can effectively support the training of LLMs for RTL code generation.
@@ 2.2. Reflection for Verilog Generation
Recognizing the limitations of previous research on LLMs for RTL generation, AutoChip  (Thakur et al., 2023c) and RTLFixer  (Tsai et al., 2023) decide to introduce self-reflection into the their framework. Specifically, RTLFixer employs GPT to generate and reflect on code. It includes approximately 80 erroneous code snippets that fail compilation, each paired with guidance. During the RTL code generation process, if the generated code fails to compile, RTLFixer  (Tsai et al., 2023) queries the database for the guidance most closely related to the error. However, the generalization of this approach is limited due to the constraint of the guidance database. In contrast, AutoChip  (Thakur et al., 2023c) directly provides the compiler error messages to the LLM rather than querying a database for guidance.
@@
2.3. Hardware Programming
Hardware description languages such as Verilog or VHDL are widely adopted for hardware design. As an alternative, High-Level Synthesis (HLS) is a crucial methodology for hardware design (Canis et al., 2011; Coussy et al., 2009). Similar to using LLMs to write RTL code, HLS enables designers to specify hardware functionality using high-level programming languages like C, C++, or SystemC. These specifications are then automatically converted into hardware description languages such as VHDL or Verilog (Xu et al., 2022; Dai and Zhang, 2019; Josipović et al., 2018; Hsiao and Anderson, 2019). By automating this translation process, HLS simplifies the design of complex hardware systems, reducing the time and effort required to manage intricate low-level coding, allowing designers to optimize and manipulate designs more efficiently (Jia et al., 2021, 2022).
@@ 3.1. Overview
The overview of the code-to-code augmentation methodology and OriGen’s generation and self-reflection framework are illustrated in Figure 1 and Figure 2, respectively.
As shown in Figure 1, our code-to-code augmentation process generates two datasets: an enhanced code dataset and an error-correction dataset. The enhanced code dataset pairs Verilog code with natural language descriptions, enabling the LLM to learn the mapping between RTL code structures and high-level semantics. The error-correction dataset contains instructions erroneous, code samples and their corrections, enhancing the LLM’s self-reflection and error correction capabilities. This dataset is constructed by collecting cases where generated code fails evaluation, documenting the specific errors and compiler feedback, and providing corrected versions of the code that pass evaluation.
As shown in Figure 2, in the code generation and error correction stage, OriGen comprises a base LLM and two trained LoRA (Low-Rank Adaptation) models: Gen LoRA and Fix LoRA. Gen LoRA is fine-tuned on the enhanced code dataset. It specializes in interpreting natural language instructions and generating initial RTL code that aligns with the given specifications. Fix LoRA is fine-tuned on the error-correction dataset and built upon Gen LoRA. This component is responsible for analyzing compiler error messages, identifying code issues, and proposing and implementing fixes to erroneous code. OriGen receives natural language instructions as input, and Gen LoRA generates initial RTL code based on these instructions. The generated code then undergoes compilation evaluation. If the code fails to pass evaluation, it enters the self-reflection loop. In this loop, Fix LoRA analyzes the error messages and the erroneous code, proposes and implements fixes to address the identified issues, and the corrected code is re-evaluated. This process iterates until the code passes verification or reaches a predefined maximum number of iterations.
@@ 3.2. Code-to-Code Augmentation
A full example of dataset augmentation is shown in Figure 4. The process begins with the original Verilog code, which defines a flip-flop with synchronous reset and enable signals. The LLM generates a description of this code, explaining its functionality, specifically, how the output q is updated based on the input d on the rising edge of the clock clk. Using this description, the LLM produces an augmented version of the code. The enhanced code presents a more concise and organized structure while maintaining the same functionality, demonstrating the effectiveness of LLMs in improving RTL code readability and quality.
@@
3.3. Error-Correction Dataset
Given the observed performance gap, we decide to adopt the similar method to acquire stronger error understanding and self-reflection capabilities from the closed-source LLM. During the code-to-code augmentation process discussed in Section 3.2, the closed-source LLM generates a diverse set of RTL code samples that fail to compile, which are then corrected in the subsequent self-reflection process. We select the code samples generated in the code-to-code augmentation process that pass compilation after correction, along with the corresponding samples that failed to compile. The natural language instruction descriptions and compiler error messages associated with these samples serve as data sources for constructing the error-correction dataset. This approach enables the model to be exposed to a broad spectrum of error types and their respective corrections, facilitating the development of robust error understanding and self-reflection capabilities.
3.4. Code Generation and Fix